# Graviscalar quasinormal modes and asymptotic tails of a thick brane

### v2

@@ Abstract
In this work, we investigate the graviscalar quasinormal modes (QNMs) and their asymptotic tail behavior of a thick brane. Considering the scalar perturbations of the thick brane metric, we obtain the main equations of graviscalar Kaluza-Klein modes. Based on these equations, the frequencies of the graviscalar QNMs of the thick brane are obtained by the Wentzel-Kramers-Brillouin, asymptotic iteration, and numerical evolution methods. The results show that the scalar fluctuation of the thick brane has a series of discrete QNMs, similar to the tensor perturbation of the brane. These modes appear as decaying massive scalar particles in four-dimensional spacetime. We also studied in detail the late time tails of these QNMs and found that some modes have slowly decaying oscillatory tails that may be new sources of the gravitational wave backgrounds. Obviously, the QNMs contain the information of the brane and are characteristic modes of the thick brane.
@@ I Introduction
Since the 20th century, the idea that our universe might be embedded in a higher-dimensional spacetime as a four-dimensional object has sparked widespread interest in the spacetime view. The early motivation behind extra-dimensional theories was to unify various interactions in nature, such as the Kaluza-Klein (KK) theory and string theory kaluza:1921un ; Klein:1926tv ; Scherk:1974ca . With the development of extra-dimensional theories, some models focusing on addressing prominent issues in physics emerged, such as the Randall-Sundrum (RS) curved extra-dimensional model Randall:1999ee ; Randall:1999vf . To explain the longstanding hierarchy problem in particle physics, namely the huge disparity between the weak scale and the Planck scale, Randall and Sundrum proposed the RS-I brane model composed of two branes embedded in a five-dimensional anti-de Sitter spacetime. Subsequently, based on the RS-I two-brane model, Randall and Sundrum further proposed the RS-II model with an infinite extra dimension. The inspiration behind the RS-II model was that even with an infinite extra dimension, the branes could still recover a four-dimensional effective gravitational potential. Due to this novel and intriguing characteristic, the RS-II model garnered widespread attention and has seen significant development in cosmology, particle physics, and black hole physics Shiromizu:1999wj ; Tanaka:2002rb ; Gregory:2008rf ; Bauer:2016lbe ; Jaman:2018ucm ; Adhikari:2020xcg ; Bhattacharya:2021jrn ; Geng:2020fxl ; Geng:2021iyq ; Geng:2022dua .
@@
In the RS-II braneworld model, our world is embedded as a four-dimensional hypersurface in a five-dimensional spacetime, with the energy density along the extra dimension distributed as a delta function, hence termed the brane model. The thick brane model, which corresponds to this, is a smooth extension of the RS-II brane model, possessing a rich internal structure and thus has richer properties DeWolfe:1999cp ; Gremm:1999pj ; Csaki:2000fc . In thick brane theories, the localization of the gravitational zero mode and various matter fields is necessary. Currently, various thick brane solutions of gravity theories and the localization of gravitational zero modes and matter fields on the brane have been studied Afonso:2007gc ; Dzhunushaliev:2010fqo ; Dzhunushaliev:2011mm ; Geng:2015kvs ; Melfo2006 ; Almeida2009 ; Zhao2010 ; Chumbes2011 ; Liu2011 ; Bazeia:2013uva ; Xie2017 ; Gu2017 ; ZhongYuan2017 ; ZhongYuan2017b ; Zhou2018 ; Hendi:2020qkk ; Xie:2021ayr ; Moreira:2021uod ; Xu:2022ori ; Silva:2022pfd ; Xu:2022gth . However, it is important to note that the localization of the zero modes of gravity and various matter fields is aimed at recovering effective physics on the four-dimensional brane, but this does not bring us new insights. To capture the extra dimension, we are more concerned with things beyond these zero modes, namely the massive KK modes that contain information about the extra dimension.
@@
It is generally believed that apart from the zero modes, thick branes also have a continuum of massive KK modes. Our recent studies indicate that within these continuous KK modes, there exists a series of discrete gravitational quasinormal modes (QNMs) Tan:2022vfe ; Tan:2023cra ; Jia:2024pdk . The spectrum of these QNMs is entirely determined by the structure of the thick brane, and conversely, these modes also contain crucial information about the extra dimension. Previous studies of the QNMs of the braneworld only considered the transverse-traceless tensor fluctuations of the metric, corresponding to spin-2 gravitons on the brane. For the RS-II thin brane model, this is sufficient because, without considering material sources on the brane, the metric fluctuation of the RS-II thin brane consists only of the transverse-traceless tensor mode Randall:1999vf . However, for the thick brane scenario, the situation is quite different. Since thick branes are generated by a background scalar field coupled with gravity, their complete metric fluctuations of thick branes should include tensor fluctuations, vector fluctuations (graviphoton), and scalar fluctuations (graviscalar) Giovannini:2001fh ; Giovannini:2001xg ; Giovannini:2001vt ; Kobayashi:2001jd . Vector fluctuations and scalar fluctuations contain information about the matter field that constitutes the thick brane, which may be the key to distinguishing between the thick brane model and the thin brane model. Additionally, in braneworld cosmology, scalar fluctuations of branes are also an important subject of study Langlois:2000iu ; Koyama:2004ap ; Hiramatsu:2004aa ; Brax:2004xh ; Maartens:2010ar ; Maier:2013gua ; Jaman:2018ucm ; Banerjee:2020uil ; Ravanpak:2022awx .
@@
For vector fluctuations, generally only non-localized zero mode exist, with no other modes. Tensor fluctuations exhibit a bound zero mode and an infinite number of massive KK modes. As for the scalar fluctuations of thick branes, in most cases, zero modes exist but cannot be bound to the brane, and the massive mode is similar to the case of tensor fluctuations. That’s all we know. So, the question now is, can we know more about the scalar fluctuations of thick branes? Do scalar fluctuations also exhibit QNMs like tensor fluctuations? By employing the methods developed from studying the QNMs of thick brane tensor fluctuations, we will investigate the QNMs of flat thick brane scalar fluctuations in this paper, aiming to understand more about scalar fluctuations.
@@ II Thick brane solution and its scalar perturbation
By using the superpotential method, the thick brane solution is given in Ref Gremm:1999pj :
@@
where $k$ is a constant with mass dimension one. Plots of the warp factor, scalar field, and scalar potential mentioned above are shown in Fig. 1.
(a)  the warp factor (8)
(b)  the scalar field (9)
(c)  the scalar potential (10)
Figure 1: The shapes of the warp factor (8), the scalar field (9), and the scalar potential (10).
@@
[…]
@@
$-\partial_{t}^{2}\tilde{\Psi}+\partial_{z}^{2}\tilde{\Psi}-U(z)\tilde{\Psi}-a^{2}\tilde{\Psi}=0,$ (22)
@@
is the effective potential and $a=\sqrt{a^{i}a_{i}}$ is a constant. Then we decompose the function $\tilde{\Psi}$ further as $\tilde{\Psi}=e^{-i\omega t}\psi(z)$ and substitute it into the above wave equation (22), a Schrödinger-like equation can be obtained as
$-\partial_{z}^{2}\psi(z)+U(z)\psi(z)=m^{2}\psi(z),$ (24)
@@
We plot the effective potential $U(z)$ in Fig. (2). We can see that the effective potential is a pure barrier and $U(\pm\infty)\rightarrow 0$. In addition, the specific form of the scalar zero mode is
@@
Obviously, this zero mode cannot be localized on the brane. Now we turn our attention to the massive KK modes. Since the effective potential (30) approaches to 00 at $kz=\pm\infty$, these scalar massive KK modes always tunnel to extra dimension infinity. Thus, the boundary conditions of these massive KK modes are
$\psi(z)\propto\left\{\begin{aligned} e^{imz},&~{}~{}~{}~{}~{}z\to\infty.&\\ e^{-imz},&~{}~{}~{}~{}~{}z\to-\infty,&\end{aligned}\right.$ (32)
@@
$WKB~{}method$ The WKB method is a very common method for solving the QNMs of a black hole Iyer:1986np . This method is based on matching the asymptotic WKB solutions at the event horizon and space infinity with the Taylor expansion near the peak of effective potential barrier through the two turning points. Since the structure of the effective potential of the scalar perturbation is similar to that of a black hole, the WKB method is also suitable for thick brane cases. Here we use the WKB method to solve the graviscalar QNMs of the thick brane. The complex frequency formula of the sixth order WKB method is Konoplya:2003ii
@@
$\;\;n\;\;$   WKB method         AIM
    $\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$ ${}~{}~{}~{}~{}~{}\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$
@@
$Asymptotic~{}iteration~{}method$ In order to accurately solve high overtone modes, we use the asymptotic iteration method (AIM). The AIM is an effective method to find the eigenvalues of linear second-order homogeneous differential equation Ciftci:2003As ; ciftci:2005co . In general, if the equation has the following form
@@
The QNFs can be solved from the above “quantization condition”. However, computing the recurrence relations (39) and (40) requires a lot of resources. Thus, Cho $et~{}al$ improved the original AIM by using the Taylor series Cho:2011sf . This greatly improved the accuracy and speed of numerical calculation. Expanding the functions $\lambda_{n}(x)$ and $s_{n}(x)$ at the point $\chi$:
@@
$\psi(u)\propto\left\{\begin{aligned} e^{-\frac{im/k}{2u-2}},&~{}~{}~{}u\to 1,&\\ e^{\frac{im/k}{2u+2}},&~{}~{}~{}u\to-1.&\end{aligned}\right.$ (49)
@@
×$\times\Bigg{[}3\left(6u^{4}-7u^{2}+6\right)\left(u^{2}+1\right)^{3}$
@@
$Time~{}evolution$ Both AIM and WKB methods study the graviscalar QNFs of the thick brane in the frequency domain. These frequencies tell us how perturbations decay in the late time. However, this lacks an understanding of the relative amplitudes of these fluctuations. Therefore, we should study the time evolution of the perturbed scalar field. Here, we use the null coordinates $u=t-z$ and $v=t+z$ to perform the time evolution. The evolution equation (22) in the null coordinates is
@@
This Gaussian pulse is located at $kv_{c}=5$ and has a width of $k\sigma=1$. We extract the data at $kz_{\text{ext}}=30$. The result of the time evolution of the Gaussian pulse is shown in Fig. 4(a). It can be seen that in logarithmic coordinates, the waveform is divided into three stages: the initial burst stage, the exponential damping stage, and the power-law tail stage. Since the QNM with the smallest imaginary part dominates the evolution, we can extract the frequency of the QNM with the smallest imaginary part by fitting the data. The fitted QNF for the case of Fig. 4(a) is $\frac{m}{k}=1.999394-0.503601i$, which is in good agreement with the result of the AIM and the WKB method. In addition to the first QNF, we can also extract the frequency of the second QNM due to the symmetry of the effective potential. When the initial data is given as a Gaussian wave packet, both odd and even QNMs are excited. The first QNM is even, and the odd QNM is covered. If the initial data is given as a static odd wave packet, only odd QNMs will be excited. In this way, we can extract the frequency of the first odd QNM. The odd initial data is given by
@@
We extract the data at $kz_{\text{ext}}=30$. The result is shown in Fig. 4(b) and the fitted QNF is $\frac{m}{k}=1.576312-1.746200i$. This is also consistent with the result of the AIM and the WKB method. For the high overtone modes, we look forward to developing more methods in the future to compare them with the results of the AIM.
@@
The above results show that a series of discrete QNMs exist in the scalar perturbation of the metric of the thick brane, just like the tensor perturbation of the thick brane. These graviscalar QNMs appear as dissipative scalar particles on the brane, corresponding to the scalar degrees of freedom of gravity. In addition, the frequencies of the first three QNMs of tensor fluctuation are Tan:2022vfe : $\frac{m_{1}}{k}=0.997018-0.526362i$, $\frac{m_{2}}{k}=0.581489-1.851280i$, and $\frac{m_{3}}{k}=0.306006-3.53366i$. Comparing the first three QNMs of scalar fluctuation and tensor fluctuation of thick brane, we can find that the real part of the first three graviscalar QNMs is larger than the tensor QNMs, but the imaginary part is smaller than the tensor QNMs. This shows that the graviscalar QNMs have a higher oscillation frequencies and a longer lifetimes. Moreover, since the original RS-II thin brane does not have scalar fluctuations, the graviscalar QNMs of the thick brane may be a key feature that distinguishes thin brane from thick brane. If such a QNM is detected, it is conclusive evidence for the existence of extra dimensions and the brane has a thickness. But because these modes have such a short lifetime (when $k=10^{-3}$eV, the lifetime of the first graviscalar QNM of the thick brane is $\tau\approx 10^{-13}$s), they are nearly impossible to detect with existing gravitational wave detectors. Nevertheless, these graviscalar QNMs may play an important role in the early universe by coupling to the 4-dimensional matter field, in particular the evolution of the early universe and the anisotropy of the cosmic microwave background radiation Seahra:2005wk ; Seahra:2005iq . On the other hand, the discussion above is for the case of $a=0$. When $a\neq 0$, these QNMs have a slowly decaying oscillation tail in the late time. Recently, R. A. Konoplya and A. Zhidenko suggests that these tails of massive patterns from extra dimensions may contribute to the gravitational wave background Konoplya:2023fmh . We expect to find signals of them in existing or future gravitational wave background detectors, such as the pulsar timing array.
(a)  Gaussian pulse
(b)  odd wave packet
Figure 4: (a): Time evolution of the Gaussian pulse at $kz=30$. (b): Time evolution of the odd wave packet at $kz=30$.
@@
Figure 5: Upper panel: Time evolution of the Gaussian pulse with different $a$ at $kz_{\text{ext}}=30$. Lower panel: Time evolution of the odd wave packet with different $a$ at $kz_{\text{ext}}=30$.
$\;\;a\;\;$   Gauss pulse         Odd wave packet
  $\alpha$    $\beta~{}~{}$ ${}~{}~{}~{}\alpha$    $\beta~{}~{}$

### v3

@@ Abstract
In this work, we investigate the quasinormal modes (QNMs) and the late time oscillatory tails of graviscalar perturbations in a thick brane. By considering the scalar perturbations of the thick brane metric, we obtain the main equations of graviscalar Kaluza-Klein modes. Using these equations, the frequencies of the graviscalar QNMs of the thick brane are obtained by the Wentzel-Kramers-Brillouin, asymptotic iteration, and numerical evolution methods. The results show that the scalar fluctuation of the thick brane has a series of discrete QNMs. In the four-dimensional spacetime, these modes manifest as decaying massive scalar particles. We also study in detail the late time tails of the scalar fluctuation. We find that the brane has slowly decaying oscillatory tails that may be new sources of the gravitational wave backgrounds. These novel phenomena may have potential phenomenological significance, and signals from these extra dimensions may potentially be found in current or future gravitational wave detectors.
@@ I Introduction
Since the 20th century, the idea that our universe might be embedded in a higher-dimensional spacetime as a four-dimensional object has sparked widespread interest in the spacetime view. The early motivation behind extra-dimensional theories was to unify various interactions in nature, such as the Kaluza-Klein (KK) theory and string theory kaluza:1921un ; Klein:1926tv ; Scherk:1974ca . With the development of extra-dimensional theories, some models focusing on addressing prominent issues in physics emerged, such as the Randall-Sundrum (RS) curved extra-dimensional model Randall:1999ee ; Randall:1999vf . To explain the longstanding hierarchy problem in particle physics, namely the huge disparity between the weak scale and the Planck scale, Randall and Sundrum proposed the RS-I brane model composed of two branes embedded in a five-dimensional anti-de Sitter spacetime. Subsequently, based on the RS-I two-brane model, Randall and Sundrum further proposed the RS-II model with an infinite extra dimension. The inspiration behind the RS-II model was that even with an infinite extra dimension, the branes could still recover a four-dimensional effective gravitational potential. Due to this novel and intriguing characteristic, the RS-II model garnered widespread attention and has seen significant development in cosmology, particle physics, and black hole physics Shiromizu:1999wj ; Tanaka:2002rb ; Zhidenko:2008fp ; Gregory:2008rf ; Konoplya:2011qq ; Bauer:2016lbe ; Jaman:2018ucm ; Adhikari:2020xcg ; Bhattacharya:2021jrn ; Geng:2020fxl ; Geng:2021iyq ; Geng:2022dua .
@@
In the RS-II braneworld model, our world is embedded as a four-dimensional hypersurface in a five-dimensional spacetime, with the energy density along the extra dimension distributed as a delta function, hence it is termed the brane model. The thick brane model, which corresponds to this, is a smooth extension of the RS-II brane model, possessing a rich internal structure and thus has richer properties DeWolfe:1999cp ; Gremm:1999pj ; Csaki:2000fc . In thick brane theories, the localization of the gravitational zero mode and various matter fields is necessary. Various thick brane solutions of gravity theories and the localization of gravitational zero modes and matter fields on the brane have been studied Afonso:2007gc ; Dzhunushaliev:2010fqo ; Dzhunushaliev:2011mm ; Geng:2015kvs ; Melfo2006 ; Almeida2009 ; Zhao2010 ; Chumbes2011 ; Liu2011 ; Bazeia:2013uva ; Xie2017 ; Gu2017 ; ZhongYuan2017 ; ZhongYuan2017b ; Zhou2018 ; Hendi:2020qkk ; Xie:2021ayr ; Moreira:2021uod ; Xu:2022ori ; Silva:2022pfd ; Xu:2022gth . However, it is important to note that the localization of the zero modes of gravity and various matter fields is aimed at recovering effective physics on the four-dimensional brane, but this does not bring us new insights. To explore the effects of the extra dimension, we are more interested in the massive KK modes beyond these zero modes, namely the massive KK modes that contain information about the extra dimension.
@@
It is generally believed that apart from the zero modes, thick branes also have a continuum of massive KK modes. Our recent studies have indicated that within these continuous KK modes, there exists a series of discrete gravitational quasinormal modes (QNMs) Tan:2022vfe ; Tan:2023cra ; Jia:2024pdk ; Tan:2024url . The spectrum of these QNMs is entirely determined by the structure of the thick brane, and conversely, these modes also contain crucial information about the extra dimension. Previous studies of the QNMs of the braneworld only considered the transverse-traceless tensor fluctuations of the metric, corresponding to spin-2 gravitons on the brane. For the RS-II thin brane model, this is sufficient because, without considering material sources on the brane, the metric fluctuation of the RS-II thin brane consists only of the transverse-traceless tensor mode Randall:1999vf . However, for the thick brane scenario, the situation is quite different. Since thick branes are generated by a background scalar field coupled with gravity, their complete metric fluctuations of thick branes should include tensor fluctuations, vector fluctuations (graviphoton), and scalar fluctuations (graviscalar) Giovannini:2001fh ; Giovannini:2001xg ; Giovannini:2001vt ; Kobayashi:2001jd . Vector fluctuations and scalar fluctuations contain information about the matter field that constitutes the thick brane, which may be the key to distinguishing between the thick brane model and the thin brane model. Additionally, scalar fluctuations have also been widely studied in other higher-dimensional spacetimes, such as scalar fluctuations of brane cosmology Langlois:2000iu ; Koyama:2004ap ; Hiramatsu:2004aa ; Brax:2004xh ; Maartens:2010ar ; Maier:2013gua ; Jaman:2018ucm ; Banerjee:2020uil ; Ravanpak:2022awx and graviscalar QNMs of higher-dimensional black holes Cardoso:2003vt ; Natario:2004jd ; Konoplya:2008ix .
@@
For vector fluctuations, generally only non-localized zero mode exist, with no other modes. Tensor fluctuations exhibit a bound zero mode and an infinite number of massive KK modes. As for the scalar fluctuations of thick branes, in most cases, zero modes exist but cannot be bound to the brane, and the massive mode is similar to the case of tensor fluctuations. So, the question now is, can we know more about the scalar fluctuations of thick branes? Do scalar fluctuations also exhibit QNMs like tensor fluctuations? By employing the methods developed from studying the QNMs of thick brane tensor fluctuations, we will investigate the QNMs of scalar fluctuations of flat thick branes in this paper, aiming to understand more about scalar fluctuations.
@@ II Thick brane solution and its scalar perturbation
By using the superpotential method, the thick brane solution is given in Ref. Gremm:1999pj :
@@
where $k$ is a constant with mass dimension one.
@@
$\text{matter}:$ $\partial_{z}^{2}\delta\phi+3\partial_{z}A\partial_{z}\delta\phi+(4\partial_{z}\Psi-\partial_{z}\varphi-6\partial_{z}\varphi)\partial_{z}\phi$ (17)
@@
$\partial_{t}^{2}\tilde{\Psi}+\partial_{z}^{2}\tilde{\Psi}-U(z)\tilde{\Psi}-a^{2}\tilde{\Psi}=0,$ (22)
@@
is the effective potential. Parameter $a=\sqrt{a^{i}a_{i}}$ corresponds to the magnitude of the spatial three-momentum of the KK particle along the brane. When $a=0$, it means that the KK particle only moves along the extra dimension and moves at the speed of light at infinity in the extra dimension. When $a>0$, the graviscalar KK mode moves both along the extra dimension and on the brane. Then we decompose the function $\tilde{\Psi}$ further as $\tilde{\Psi}=e^{-i\omega t}\psi(z)$ and substitute it into the above wave equation (22), a Schrödinger-like equation can be obtained as
$\partial_{z}^{2}\psi(z)+U(z)\psi(z)=m^{2}\psi(z),$ (24)
@@
We plot the effective potential $U(z)$ in Fig. 1. We can see that the effective potential is a pure barrier and $U(\pm\infty)\rightarrow 0$. In addition, the specific form of the scalar zero mode is
@@
Obviously, this zero mode cannot be localized on the brane. Now we turn our attention to the massive KK modes. Since the effective potential (30) approaches to 0 at $kz=\pm\infty$, these scalar massive KK modes always tunnel to extra dimension infinity. Thus, the boundary conditions of these massive KK modes are
[…]
@@
WKBmethodWKB~method The WKB method is a very common method for solving the QNMs of a black hole Iyer:1986np ; Kanti:2006ua . This method is based on matching the asymptotic WKB solutions at the event horizon and space infinity with the Taylor expansion near the peak of effective potential barrier through the two turning points. The complex frequency formula of the sixth order WKB method is Konoplya:2003ii
@@
n\;\;n\;\;   WKB method         AIM
    $\text{Re}(m/k)$    $\text{Im}(m/k)$~~ $\text{Re}(m/k)$    $\text{Im}(m/k)$~~
@@
AsymptoticiterationmethodAsymptotic~iteration~method In order to accurately solve high overtone modes, we use the asymptotic iteration method (AIM). The AIM is an effective method to find the eigenvalues of linear second-order homogeneous differential equation Ciftci:2003As ; ciftci:2005co . In general, if the equation has the following form
@@
The QNFs can be solved from the above “quantization condition”. However, computing the recurrence relations (39) and (40) requires a lot of resources. Thus, Cho etalet~al improved the original AIM by using the Taylor series Cho:2011sf . This greatly improved the accuracy and speed of numerical calculation. Expanding the functions $\lambda_{n}(x)$ and $s_{n}(x)$ at the point $\chi$:
@@
[…]
@@
$\times\Bigg[3\left(6u^{4}-7u^{2}+6\right)\left(u^{2}+1\right)^{3}$
@@
The above results are obtained by choosing $a=0$, but since $\omega^{2}=a^{2}+m^{2}$, we can easily know the influence of the separation variable parameter a on the QNF: the real part of $\omega$ increases with $a$, while the imaginary part of $\omega$ decreases with $a$. This means that both the oscillation frequency and lifetime of the graviscalar QNMs on the brane increase with $a$. In addition, the value of $a$ will also significantly affect the tail behavior of the graviscalar QNMs, which will be seen in the next section.
@@
TimeevolutionTime~evolution Both AIM and WKB methods study the graviscalar QNFs of the thick brane in the frequency domain. These frequencies tell us how perturbations decay in the late time. However, this lacks an understanding of the relative amplitudes of these fluctuations. Therefore, we should study the time evolution of the perturbed scalar field. Here, we use the null coordinates u=tzu=t-z and $v=t+z$ to perform the time evolution. The evolution equation (22) in the null coordinates is
@@
This Gaussian pulse is located at $kv_{c}=5$ and has a width of $k\sigma=1$. We extract the data at $kz_{\text{ext}}=10$. The result of the time evolution of the Gaussian pulse is shown in Fig. 3(a). It can be seen that in logarithmic coordinates, the waveform is divided into three stages: the initial burst stage, the exponential damping stage, and the power-law tail stage. Since the QNM with the smallest imaginary part dominates the evolution, we can extract the frequency of the QNM with the smallest imaginary part by fitting the data. The fitted QNF for the case of Fig. 3(a) is $\frac{m}{k}=1.999394-0.503601i$, which is in good agreement with the result of the AIM and the WKB method. In addition to the first QNF, we can also extract the frequency of the second QNM due to the symmetry of the effective potential. When the initial data is given as a Gaussian wave packet, both odd and even QNMs are excited. The first QNM is even, and the odd QNM is covered. If the initial data is given as a static odd wave packet, only odd QNMs will be excited. In this way, we can extract the frequency of the first odd QNM. The odd initial data is given by
@@
We extract the data at $kz_{\text{ext}}=10$. The result is shown in Fig. 3(b) and the fitted QNF is $\frac{m}{k}=1.576312-1.746200i$. This is also consistent with the result of the AIM and the WKB method. For the high overtone modes, we look forward to developing more methods in the future to compare them with the results of the AIM.
@@
$U(z)\sim\frac{\frac{3}{2}(\frac{3}{2}+1)k^{2}}{k^{2}z^{2}}.$ (63)
(a)  Gaussian pulse
(b)  odd wave packet
Figure 3: (a): Time evolution of the Gaussian pulse at $kz=10$. (b): Time evolution of the odd wave packet at $kz=10$.
@@
Figure 4: Upper panel: Time evolution of the Gaussian pulse with different $a$ at $kz_{\text{ext}}10$. Lower panel: Time evolution of the odd wave packet with different $a$ at $kz_{\text{ext}}=10$.
a\;\;a\;\;   Gauss pulse         Odd wave packet
  $\alpha$    $\beta$~~ $\alpha$    $\beta$~~
@@
The above results show that a series of discrete QNMs exist in the scalar perturbation of the metric of the thick brane, just like the tensor perturbation of the thick brane. These graviscalar QNMs appear as dissipative scalar particles on the brane, corresponding to the scalar degrees of freedom of gravity. In addition, the frequencies of the first QNM of tensor fluctuation is Tan:2022vfe : $\frac{m_{1}}{k}=0.997018-0.526362i$. Comparing the first QNM of scalar fluctuation and tensor fluctuation of thick brane, we can find that the real part of the first graviscalar QNM is larger than the tensor QNM, but the imaginary part is smaller than the tensor QNM. This shows that the graviscalar QNMs have a higher oscillation frequencies (the oscillation frequency of the first graviscalar is twice that of the corresponding tensor mode) and a slightly longer lifetime. Moreover, since the original RS-II thin brane does not have scalar fluctuations, the graviscalar QNMs of the thick brane may be a key feature that distinguishes thin brane from thick brane. If such a QNM is detected, it is conclusive evidence for the existence of extra dimensions and the brane has a thickness. However, due to the short lifetime and extremely high frequency of these modes (when $k=10^{-3}$ eV and $a/k=0$, the lifetime of the first QNM of the thick brane is $\tau\approx 10^{-13}$ s, and the oscillation frequency is about $10^{23}$ Hz), although the lifetime increases with $a$, the oscillation frequency also increases. Therefore, it is almost impossible to detect them with existing gravitational wave detectors. Perhaps high-frequency gravitational wave detectors under construction, such as The CERN Axion Solar Telescope Ejlli:2019bqj , could capture traces of them. Nevertheless, these graviscalar QNMs may play an important role in the early universe by coupling to the 4-dimensional matter field, in particular the evolution of the early universe and the anisotropy of the cosmic microwave background radiation Seahra:2005wk ; Seahra:2005iq . On the other hand, when $a\neq 0$, these QNMs have a slowly decaying oscillation tail in the late time. Recently, R. A. Konoplya and A. Zhidenko suggests that these tails of massive patterns from extra dimensions may contribute to the gravitational wave background Konoplya:2023fmh . We expect to find signals of them in existing or future gravitational wave background detectors, such as the pulsar timing array.